OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) [02.03] Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No solution Question 2 (Multiple Choice Worth 2 points) [02.03] The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution Question 3 (Multiple Choice Worth 2 points) [02.03] Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution Question 4 (Multiple Choice Worth 2 points) [02.03] Solve for x: −3|x + 7| = −12 x = 5 over 3, x = −19 over 3 x = −3, x = −11 x = −3, x = 11 No solution Question 5 (Multiple Choice Worth 2 points) [02.03] Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x − 6| = −41 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined. Question 6 (Multiple Choice Worth 1 points) [02.03] Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution

OpenStudy (anonymous):

Whoever tells only answers will get Fan and Medal

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

thats confusing af

OpenStudy (anonymous):

:/

OpenStudy (jiteshmeghwal9):

What's so confusing ?

OpenStudy (anonymous):

The forumula you gave me

OpenStudy (anonymous):

Formula*

OpenStudy (anonymous):

actually

OpenStudy (anonymous):

The only questions I can't get are

OpenStudy (jiteshmeghwal9):

ok let x=5 |5|=5 here x=5>0 |-5|=-(-5)=5 because x=-5<0

OpenStudy (anonymous):

The ones that are like: |x − 2| + 10 = 12

OpenStudy (anonymous):

What would I do for that?

OpenStudy (jiteshmeghwal9):

the mod gives us absolute value & its value always comes positive.

OpenStudy (jiteshmeghwal9):

you are asked here to find a value for x which satisfy the eqn.

OpenStudy (anonymous):

I mean is there like a formula for that? cause I get the ones that are like 2|x+7| = 4 or whatever

OpenStudy (anonymous):

I don't get the ones where you add

OpenStudy (anonymous):

Also the ones like 3|x − 8| + 15 = 18

OpenStudy (jiteshmeghwal9):

\[|x-2|+10=12\]\[|x-2|=2\]\[(x-2)=2.....eqn1\]\[-(x-2)=2...eqn2\]i'm sorry for the silly mistake in typing

OpenStudy (anonymous):

Ahhhhhh that makes sense!!!

OpenStudy (anonymous):

You just earned a medal and a fan :D still might need help to make sure I got the right answers

OpenStudy (anonymous):

For 1 is it No Solution?

OpenStudy (jiteshmeghwal9):

\[|x|=x \] for x>0\[|x|=-x\]for x<0 this is ur formula

OpenStudy (jiteshmeghwal9):

no, the answer is option 1 for ur 1st question

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

For the second one is it No Solution? Cuz I got x=9 and x=8

OpenStudy (anonymous):

?

OpenStudy (jiteshmeghwal9):

no i think u r doing a mistake becuz the answer is x=9 & 7

OpenStudy (anonymous):

hmm

OpenStudy (jiteshmeghwal9):

|x-8|=1 -(x-8)=1 (x-8)=1

OpenStudy (jiteshmeghwal9):

these two eqns will give u the answer for 2nd question

OpenStudy (anonymous):

Ungh I only got like 2 right

OpenStudy (anonymous):

gtg now

OpenStudy (anonymous):

Guys

OpenStudy (anonymous):

Whoever answers the questions gets medal and fan

OpenStudy (anonymous):

Only tell answers plz

OpenStudy (anonymous):

1)A 2)C 3)B 4)B 5)B 6)D I tried, sorry if some are wrong.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

Thank you SO OOOO much

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

emm I have another small quiz

OpenStudy (anonymous):

can you do that? I already fanned I'll give medal if you can do that

OpenStudy (anonymous):

Sure!!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) [02.03] Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution Question 2 (Multiple Choice Worth 2 points) [02.03] Solve for x: |x| − 8 = −5 x = −13 and x = −3 x = 3 and x = −3 x = 3 and x = 13 No solution Question 3 (Multiple Choice Worth 2 points) [02.03] Compare and Contrast: Two equations are listed below. Solve each equation and compare the solutions. Choose the statement that is true about both solutions. Equation 1 Equation 2 |5x − 6| = −41 |7x + 13| = 27 Equation 1 has more solutions than equation 2. Equation 1 and Equation 2 have the same number of solutions. Equation 2 has more solutions than Equation 1. The number of solutions cannot be determined. Question 4 (Multiple Choice Worth 2 points) [02.03] The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution Question 5 (Multiple Choice Worth 2 points) [02.03] Solve for x: −4|x + 5| = −16 x = 21 over 4, x = −11 over 4 x = −1, x = 9 x = −1, x = −9 No solution Question 6 (Multiple Choice Worth 1 points) [02.03] Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (anonymous):

Ty so much I need this quiz done it'll help me so much

OpenStudy (anonymous):

I think these are the same questions.

OpenStudy (anonymous):

they're different

OpenStudy (anonymous):

1)C 2)B 3)C 4)C 5)C 6)B

OpenStudy (anonymous):

I am so thankful :D

OpenStudy (anonymous):

OMG